OpenStudy (anonymous):

10x - 3y=60 the x and y do not intercept

OpenStudy (anonymous):

What are you solving for?

OpenStudy (anonymous):

i dont even know thats why i posted

OpenStudy (anonymous):

Umm, well then idk. You could solve for y or x, idk which would be right. Or maybe its asking for both y and x.

OpenStudy (anonymous):

Sorry, im not sure.

OpenStudy (anonymous):

yeah im stumped

OpenStudy (anonymous):

Maybe it's asking to have the equation solved for the y-intercept form. Are there instructions for this question?

OpenStudy (shamim):

\[10x-3y=60\]

OpenStudy (shamim):

\[or, 10x=3y+60\]

OpenStudy (shamim):

\[or, x=\frac{ 3 }{ 10 }y+6\]

OpenStudy (shamim):

is it ur result which u wanted

OpenStudy (anonymous):

no, no instructions

OpenStudy (anonymous):

it just says the x and y do not intercept

OpenStudy (anonymous):

10x - 3y=60 the x and y do not intercept there is no solution then :P